OpenStudy (anonymous):

HELP! Pic Below I GIVE MEDALS ;) solve for x y z round values to the whole number

OpenStudy (anonymous):



OpenStudy (ryan123345):

I sorta stole your other question so you can do this one if you want :) @Eyad

OpenStudy (anonymous):

just use rules of S=o/h C=a/h T=o/a

OpenStudy (anonymous):

@ryan123345 :No i'am good ,Continue :P

OpenStudy (anonymous):

i don't kno that rule

OpenStudy (ryan123345):

okay :)

OpenStudy (ryan123345):

\[x=(\frac{ 28 }{ 45 })\tan^{-1} \]

OpenStudy (ryan123345):

thats for your first triangle :)

OpenStudy (ryan123345):

\[z=(\frac{ 77 }{ 85 })\sin^{-1} \]

OpenStudy (ryan123345):

Thats for your second triangle :)

OpenStudy (anonymous):

0.018=x y=0.016 z=64.94

OpenStudy (ryan123345):

where is y? lol

OpenStudy (anonymous):

in the corner or the 2nd triangle

OpenStudy (ryan123345):